100 times I draw three cards from 13 to find the experimental probability of no face cards. I get .39 where as I calculated it should be .4196... How do I get the standard deviation and what does this tell me about why my experimental answer differs from my calculated?
Answer by stanbon(75887)

100 times I draw three cards from 13 to find the experimental probability of no face cards. I get .39 where as I calculated it should be .4196... How do I get the standard deviation and what does this tell me about why my experimental answer differs from my calculated?
---
Std = sqrt(0.4196*0.5804/100) = 0.04935
z(0.39) = (0.39-0.4196)/sqrt(0.4196*0.5804/100) = -0.6000
------
Your experimental result is 0.6 standard deviations below the calculated mean.
=========================================
Cheers,
Stan H.